Please make sure to explain, make sure answer is correct and neat. Thank you!

3. Show that the time taken by a particle moving along a curve y = y(x) with velocity, ds dt center on the x-axis. Hint: t = Jo x x from the point (0,0) to the point (1,1) is minimum if the curve is a circle having its 1 ds
Answer #1

ах y-y(x) dS 0 (0,0)Let the particle moves from 'O' to 'P' along the path y=y(x)

We consider a infinitesimal segment ds on the path y(x).

Futher , we assume the particle's velocity is constant while travelling the path ds.

From geometry we find ds= \sqrt{dx^2+dy^2}= [ \sqrt{1+(\frac{dy}{dx})^2}]dx=dx \sqrt{1+y'^2}

Where y'= \frac{dy}{dx}

Given the particle's velocity v=\frac{ds}{dt}=x

So the time taken by the particle to travel the distance ds is given by dt=\frac{ds}{v}= \frac{\sqrt{[1+y'^2}]dx}{x}

So the total time taken by the particle to travle from 'O' to 'P' along the path y=y(x) is given by

t= \int_{O}^{P}dt=\int_{O}^{P} \frac{ \sqrt{[1+y'^2}]dx}{x}

Or

t= \int_{O}^{P}fdx [ where f=\frac{ \sqrt{1+y'^2}}{x} ]

Now from Least Action principle, we know that 't' will be minimum if it satisfies  Euler-Lagrange differential equation i.e

\frac{\partial f}{\partial y}-\frac{d}{dx}(\frac{\partial f}{\partial y'})=0...........................(i0

Now we find

\frac{\partial f}{\partial y}=0

So equation (i) reduces to

\frac{d}{dx}(\frac{\partial f}{\partial y'})=0

Therefore we can write

\frac{\partial f}{\partial y'}= K, [ Where K is a constant]

Differentiating f with respect to y' , we get

\frac{\partial f}{\partial y'}=\frac{y'}{x \sqrt{1+y'^2}}

Hence we can write

\frac{y'}{x \sqrt{1+y'^2}}=K

Or

\frac{y'^2}{x^2 (1+y'^2)}=K^2

Or

y'^2=K^2x^2 (1+y'^2)

Or

y'^2(1-K^2x^2) =K^2x^2

Or

y'^2 =\frac{K^2x^2}{(1-K^2x^2)}

Or

y' =\frac{Kx}{\sqrt{(1-K^2x^2)}}

Or

\frac{dy}{dx} =\frac{Kx}{\sqrt{(1-K^2x^2)}}

or

dy =\frac{Kxdx}{\sqrt{(1-K^2x^2)}}

Integrating both sides we get

\int dy =\int \frac{Kxdx}{\sqrt{(1-K^2x^2)}}

Let 1-K^2x^2=v^2\Rightarrow -K^2xdx=vdv

So we get

y=-\frac{1}{K}\int \frac{vdv}{v}= -\frac{v}{K}+L [ 'L' is a constant ]

Or

y=-\frac{\sqrt{1-K^2x^2}}{K}+L

Squaring both sides we get

K^2(y-L)^2=1-K^2x^2

Or

K^2[(y-L)^2+x^2]=1.................................

Or

[(y-L)^2+x^2]=P.................[ where P=1/K^2

Now From the given points we find when x=0 , y=0

L=P

And when x=1, y=1

L=2, So K=2

So the equation of the trajectory is

x^2+(y-2)^2=2

So the time taken by the particle is minimum if the curve is a circle and the center lies on y axis